substitution
4x+y=15
-x-3y=-12

Answer by robertb(5830)

The bottom equation gives 12-3y = x. substitute this into the top equation:
4(12-3y)+y = 15,
48-12y + y = 15,
48-11y = 15,
-11y = 15-48,
-11y = -33,
y = 3, and x = 12-3*3 = 12-9 = 3.